OpenStudy (abbycross167):

Can someone please check the answer to my question? Please? 54. A student scored an 88, 92, and 76 on three math tests. What does the student need to get on the fourth test to have an average of 85 for all four tests? Joseph wants to use algebra to solve this problem. Which equation should he use? a. 4x = 256/85 b. 256 + x = 85 4 c. 85 + x = 256 4 d. 256 = 85x 3 4

OpenStudy (abbycross167):

@Directrix

OpenStudy (kittiwitti1):

Average means add everything up and divide by the amount of values available...\[\frac{88+92+76+\text{ fourth test score}}{4}=85\]

OpenStudy (kittiwitti1):

Let's make the fourth test score x. The other three values add up to 256...\[\frac{256+x}{4}=85\]Multiply both sides by 4\[4\left(\frac{256+x}{4}\right)=(85)4\rightarrow 256+x=340\]Subtract 256 from both sides\[(256+x)-256=340-256\rightarrow x=...?\]

OpenStudy (kittiwitti1):

Oh. They just want the equation... WELL. Lol I solved it for you. It's B xD

OpenStudy (kittiwitti1):

You get why it's B though, right?

OpenStudy (abbycross167):

Haha THANK YOU VERY MUCH!!

OpenStudy (abbycross167):

and yes ma'am/ sir

OpenStudy (kittiwitti1):

oye, soy chica ♀ lol